OpenStudy (anonymous):

can someone explain w^2 +15W -250 = 0

OpenStudy (turingtest):

explain what, how to factor it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how to factor? how to find the value of w?

OpenStudy (anonymous):

yes, need to find value of W

OpenStudy (turingtest):

factoring is the way to go here we need two numbers that multiply to -250 and add to 15 can you think of what those two numbers must be?

OpenStudy (anonymous):

I was trying to follow the steps of x=(-b±√(b^2-4ac))/2a but then got lost. I got as far as x=(-15±√(15-4(2)(-250) ))/2(2) but don't know if this is correct way to figure it out.

OpenStudy (anonymous):

oh, so i need to use F.O.I.L. and noth the other way?

OpenStudy (turingtest):

no, just try to factor it the quadratic formula is not necessary here read my last post

OpenStudy (anonymous):

\[w ^{2}+15w-250=0\]\[(w+25)(w-10)=0\]

OpenStudy (anonymous):

duh, i feel so silly...

OpenStudy (anonymous):

I think i'm confusing myself or thinking too hard

OpenStudy (anonymous):

then equate to zero; w+25=0 w-10=0 w=-25 w=10

OpenStudy (turingtest):

thank you jerwyn, but I was hoping to have the asker figure it out :/ @yv173 it happens to me all the time. Look for the easiest way always. It saves much stress.

OpenStudy (anonymous):

Thank you both very much, you just saved me a night of having to stay up for hours trying to figure it out.

OpenStudy (anonymous):

if the given equation is computable using factoring method, go for it. but if not, you can proceed to quadratic formula. .

OpenStudy (anonymous):

ok, thanks now I understand

OpenStudy (anonymous):

thanks turing test for the advice.

OpenStudy (turingtest):

and in case you give up on factoring try wolfram alpha: http://www.wolframalpha.com/input/?i=factor+w%5E2%2B15w%E2%88%92250 only in emergencies though, don't rely on it. The human mind is superior!

OpenStudy (anonymous):

wow, great advice, thanks again!!!